OpenStudy (pokemon23):

Name each expression based on its degree and number of terms. 75x+5 w^2=3 -5n^4=10n-10

OpenStudy (pokemon23):

someone explain it to me ;)?

OpenStudy (anonymous):

monomial binomial polynomial

OpenStudy (pokemon23):

1=monomial?

OpenStudy (anonymous):

i believe so

OpenStudy (pokemon23):

well 75x?

OpenStudy (cornzyblack):

i would have said (tha 1st one is a linear, and the second is a quadratic and the third idk).. Sorry

OpenStudy (pokemon23):

any suggestion I have to finished my math hw please lilg you got to saved my life

OpenStudy (pokemon23):

NOOOOOOOOOOOOOOOOOOOOOOOOOOOOOOOOO

OpenStudy (anonymous):

x^y is the basic form, y determine what degree it is, so 75x is like 75x^1 (1st degree) (linear function) w^2 has a x^y (y =2) (2nd degree) (quadratic function) -5n^4 (y=4) (4th degree) (quartic function)

OpenStudy (pokemon23):

WILLING TO HELP ME IN ALGEBRA 1

OpenStudy (anonymous):

ask away

OpenStudy (pokemon23):

yes :)

OpenStudy (pokemon23):

write each polynomial in standard from. Then name each polynomial based on its degree and umber of terms.

OpenStudy (pokemon23):

4x-3x^2 k^2-5+8k

OpenStudy (anonymous):

4x-3x^2 = 0 ? k^2-5+8k = 0? for both?

OpenStudy (pokemon23):

no their diffrent

OpenStudy (pokemon23):

1)4x-3x^2 =

OpenStudy (pokemon23):

2)k^2-5+8k

OpenStudy (pokemon23):

both different questions

OpenStudy (anonymous):

-3x^2 + 4x (is the first question equal to something?) K^2 + 8k = 5

OpenStudy (pokemon23):

its not equal to anything

OpenStudy (anonymous):

yeah then its what I wrote

OpenStudy (pokemon23):

1)4x-3x^2 2)k^2-5+8k 3)9z-11z^2+5z+5

OpenStudy (pokemon23):

let post it up it making confuse

OpenStudy (anonymous):

for putting them in stand form you need to follow this rules highest degree term comes first + next highest degree + next highest degree + .... = zero degree term

OpenStudy (pokemon23):

ok :)

OpenStudy (anonymous):

so for 1) Highest degree is -3x^2 (degree 2) then 4x (degree 1) so it becomes -3x^2 + 4x for 2) its K^2 (2 degree), 8k (1 degree) and -5 (0 degree) which would make K^2 + 8k -5 Try number 3 by yourself, remember highest degree term comes first, then next highest 3) 9z-11z^2+5z+5 (hint highest degree is 2 from -11z^2)

OpenStudy (pokemon23):

ya because the rest are 1

OpenStudy (anonymous):

5 is zero degrees (its like 5x^0) but x^0 = 1 so, 5*1 = 5

OpenStudy (pokemon23):

so to find the degree of -16 will it be 0?

OpenStudy (anonymous):

yup -16 = -16x^0 which is -16 *1 = -16

OpenStudy (pokemon23):

so for number 3

OpenStudy (anonymous):

basically all numbers without (x) or variables are degree 0

OpenStudy (pokemon23):

-11z^2-9z+5z+5

OpenStudy (pokemon23):

is that correct?

OpenStudy (anonymous):

yup but what is -9z + 5z =?

OpenStudy (pokemon23):

9z-11z^2+5z+5

OpenStudy (pokemon23):

-9Z+5Z+5

OpenStudy (anonymous):

-11z^2-9z+5z+5 = -11z^2-4z+5

OpenStudy (pokemon23):

you subtract the 9+5

OpenStudy (pokemon23):

didn't know

OpenStudy (anonymous):

yup because you can, combine common terms when you can

OpenStudy (pokemon23):

ok I got more questions :'(

OpenStudy (anonymous):

I got time for another round then I'll have to go sorry

OpenStudy (pokemon23):

ok

OpenStudy (anonymous):

whats the question?

OpenStudy (pokemon23):

-(2.57)^0

OpenStudy (anonymous):

what degree is -(2.57)^0 or what is -2.57^0? -2.57^0 = 1 No matter what number to the power of 0 is always 1, (except for 0)

OpenStudy (pokemon23):

4^-2

OpenStudy (anonymous):

X^-Y = 1/(X^Y) so 4^-2 = 1/(4^2)

OpenStudy (pokemon23):

1/16

OpenStudy (anonymous):

yup

OpenStudy (pokemon23):

(-5)^-2

OpenStudy (pokemon23):

78^-1

OpenStudy (anonymous):

try these out (x^-y) = 1/(x^y)

OpenStudy (pokemon23):

Simplify each expression 1/x^-7

OpenStudy (anonymous):

(-5)^-2 = 1/(-5^2) 78^-1 = 1/(78^1) 1/(X^-7) = 1/(1/X^7)

OpenStudy (pokemon23):

Simplify each expression 6a^-1c^-3/d^0

OpenStudy (anonymous):

6a^-1 * c^-3/d^0 = 1/(6a*c^3) Sorry I have to go now

OpenStudy (pokemon23):

ok thanks